# Fairness through Difference Awareness: Measuring Desired Group Discrimination in LLMs

### v1

@@ Abstract
Algorithmic fairness has conventionally adopted a perspective of racial color-blindness (i.e., difference unaware treatment). We contend that in a range of important settings, group difference awareness matters. For example, differentiating between groups may be necessary in legal contexts (e.g., the U.S. compulsory draft applies to men but not women) and harm assessments (e.g., calling a girl a terrorist may be less harmful than calling a Muslim person one). In our work we first introduce an important distinction between descriptive (fact-based), normative (value-based), and correlation (association-based) benchmarks. This distinction is significant because each category requires distinct interpretation and mitigation tailored to its specific characteristics. Then, we present a benchmark suite composed of eight different scenarios for a total of 16k questions that enables us to assess difference awareness. Finally, we show results across ten models that demonstrate difference awareness is a distinct dimension of fairness where existing bias mitigation strategies may backfire.
@@
Angelina Wang Stanford University Michelle Phan Stanford University Daniel E. Ho* Stanford University Sanmi Koyejo* Stanford University
@@ 1 Introduction
Google Gemini’s racially diverse Nazis spotlighted a structural problem in fair generative AI Robertson (2024). Other symptoms of this problem: Claude responds that military fitness requirements are the same for men and women. Gemini recommends Benedict Cumberbatch to be cast as the last emperor of China.111Details in Appendix E.1. At their core, these issues stem from a failure to distinguish between fair differentiation and harmful prejudice.
@@
We build a benchmark suite composed of eight benchmarks spanning both descriptive and normative forms.222https://github.com/Angelina-Wang/difference_awareness Each of the eight benchmarks comprises 2,000 questions, totaling 16,000 questions. Overall, we argue that Difference Awareness and Contextual Awareness are important notions of fairness that have been neglected by existing work.
@@ 2 Prior Work
Definitions of bias. The issue of “bias” being poorly defined has been highlighted in prior work Blodgett et al. (2020), emphasizing the need to connect “bias” to specific harms Goldfarb-Tarrant et al. (2023). When left unspecified, fairness is often implicitly framed as difference unaware treatment, which parallels the problem of poorly conceptualized and operationalized notions of stereotypes Blodgett et al. (2021).
@@ 6 Discussion
We also point towards promising directions for improving on difference awareness in our different settings. For descriptive tasks, retrieval-augmented generation (RAG) Lewis et al. (2020) is a popular technique to better ground responses in fact, and can prove a fruitful direction. For normative tasks, our experiments show that prompts can steer models to be more or less DiffAware. While our preliminary experiments are not promising for CtxtAware, we point towards directions, potentially using chain-of-thought Wei et al. (2022b), that encode more human input about when difference awareness should be employed. This could be at a more general level (e.g., when different treatment is required to correct for historical disparities) or more context-specific level (e.g., that only occupations with existing representation disparities for marginalized groups should have affirmative action). And finally, for correlation tasks which may naturally appear in real-world use (e.g., creative story-telling about characters), we follow the recommendation of prior works to design human-centered interventions Yee et al. (2021); Bennett et al. (2021). In other applications, e.g., when translating the gender-neutral phrase “o bir doktor” in Turkish into either “he is a doctor” or “she is a doctor” in English,666https://blog.google/products/translate/reducing-gender-bias-google-translate/ this has looked like providing multiple options to the user. This intervention pushes the user to make explicit decisions, rather than implicitly prioritizing certain choices over others and making potentially biased decisions.
@@
First, a key limitation of our benchmark suite is that, like most benchmark suites, it primarily measures upstream performance with uncertain predictability of downstream performance Wagstaff (2012). While our benchmarks are more downstream than correlation evaluations, they may still be distinct from a specific application, e.g., writing a recommendation letter Wan et al. (2023), autocompleting emails. Our intention is that performance on our benchmark suite is indicative of performance on other downstream applications. However for these reasons, our benchmark is intended to be understood on a relative rather than absolute scale. In App. E.5 we do an analysis of the within-benchmark correlation to try and better understand this, with the idea that if performance on our benchmarks have correlation with themselves, then performance on them is likely to correlate with other downstream applications which may warrant difference awareness. This is related to the problem that our benchmark is composed of multiple choice questions, which have been shown to not necessarily correlate with other kinds of uses Röttger et al. (2024a); Tam et al. (2024). However, there are benefits to multiple choice questions. Beyond being easier to analyze, there is a lower computational cost compared to open-ended responses.
@@
Second, a related limitation has to do with scope. Four of our eight benchmarks are explicitly grounded in the United States context, and while the other four may generalize to other contexts, are likely still based on Western norms and values Sambasivan et al. (2021). Additionally, examples of scenarios not included in the coverage of our benchmark suite include reclamations of slurs (i.e., members of certain identity groups using words that would otherwise be deemed inappropriate) Jeshion (2020), what composes a hate crime, additional diversity initiatives beyond affirmative action in the occupation setting, medical reasons to treat people from different demographic groups differently.
@@
Finally, we do not disaggregate our benchmarks by demographic group. In other words, the scores combine outputs on questions asking about racial differences, gender differences, and more. While in many cases it can be important to disaggregate by demographic axis, our focus in this work is on demonstrating the erasure of difference awareness as an important concept for fairness. We encourage future work to explore whether and how difference awareness varies across demographic axes.
@@ Acknowledgments
SK acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI and Google. We are appreciative of feedback from RegLab and STAIR Lab.
@@ Appendix B Literature Review
In Tbl. 4 we list the 37 existing language model fairness benchmark papers that we review, and how we categorized them for Tbl. 1 the main text. Certain benchmarks like HarmBench were left out because there was insufficient information from the paper to determine whether a difference unaware model would do well on them.
@@ Appendix E Other empirical results
The ten models we run our experiments on are Llama-3.1 8b and 70b Grattafiori et al. (2024), Mistral-0.3 7b Jiang et al. (2023), Mistral NeMo 12b Mistral AI (2024), Gemma-2 9b and 27b Team et al. (2024), GPT-4o regular and mini OpenAI et al. (2024), Claude-3.5 Sonnet and Haiku Anthropic (2024). Here we describe the results of a series of additional experiments that we run.
@@ E.1 Introduction Examples
In Sec. 1 we include a few anecdotal examples of difference unawareness. These came from the chat interfaces of Claude and Gemini in January 2025. These were conducted in order to sanity check that we could naturally solicit compelling examples of difference unawareness even without the constraints of forced choice responses Röttger et al. (2024a); Tam et al. (2024). The full prompts and responses are included in Fig. 5.

### v2

@@ Abstract
Algorithmic fairness has conventionally adopted the mathematically convenient perspective of racial color-blindness (i.e., difference unaware treatment). However, we contend that in a range of important settings, group difference awareness matters. For example, differentiating between groups may be necessary in legal contexts (e.g., the U.S. compulsory draft applies to men but not women) and harm assessments (e.g., referring to girls as “terrorists” may be less harmful than referring to Muslim people as such). Thus, in contrast to most fairness work, we study fairness through the perspective of treating people differently — when it is contextually appropriate to. We first introduce an important distinction between descriptive (fact-based), normative (value-based), and correlation (association-based) benchmarks. This distinction is significant because each category requires separate interpretation and mitigation tailored to its specific characteristics. Then, we present a benchmark suite composed of eight different scenarios for a total of 16k questions that enables us to assess difference awareness. Finally, we show results across ten models that demonstrate difference awareness is a distinct dimension to fairness where existing bias mitigation strategies may backfire.
@@
Angelina Wang Stanford University, Cornell Tech Michelle Phan Stanford University
Daniel E. Ho* Stanford University Sanmi Koyejo* Stanford University
@@ 1 Introduction
Google Gemini’s racially diverse Nazis spotlighted a structural problem in fair generative AI Robertson (2024). Other symptoms of this problem: Claude (incorrectly) responds that U.S. military fitness requirements are the same for men and women. Gemini recommends Benedict Cumberbatch to be cast as the last emperor of China.111These examples are from open-ended questions asked to the Claude and Gemini chat interfaces. Details in Appendix E.1. At their core, these issues stem from a failure to distinguish between fair differentiation and harmful prejudice.
@@
We build a benchmark suite composed of eight benchmarks spanning both descriptive and normative forms.222Anonymized dataset submitted separately. Each of the eight benchmarks comprises 2,000 questions, totaling 16,000 questions. Overall, we argue that Difference Awareness and Contextual Awareness are important notions to fairness that have been neglected by existing work.
@@ 2 Prior Work
Definitions of bias. The issue of “bias” being poorly defined has been highlighted in prior work Blodgett et al. (2020), emphasizing the need to connect “bias” to specific harms Goldfarb-Tarrant et al. (2023). When left unspecified, fairness is often implicitly framed as difference unaware treatment, which parallels the problem of poorly conceptualized and operationalized notions of stereotypes Blodgett et al. (2021). Causal fairness has considered fairness through the lens of causal pathways and counterfactuals Kusner et al. (2017); Li et al. (2025). We do not look directly at these methods, but overall argue that for a variety of scenarios, the causal pathways from demographic group to output should actually exist.
@@ 6 Discussion
We also point towards promising directions for improving on difference awareness in our different settings. While there have been many proposed distinctions in fairness evaluations (e.g., implicit versus explicit Bai et al. (2024), intrinsic vs extrinsic Cao et al. (2022)), we believe our differentiation of descriptive, normative, and correlation to be valuable for creating targeted interventions, as each category warrants distinct treatment. For descriptive tasks, a fruitful direction is retrieval-augmented generation (RAG) Lewis et al. (2020), which is a popular technique to better ground responses in fact. For normative tasks, our experiments show that prompts can steer models to be more or less DiffAware. While our preliminary experiments are not promising for CtxtAware, we point towards directions such as further systematizing the concept of “fairness” to encode more human input Wallach et al. (2025), or using chain-of-thought Wei et al. (2022b) and other reasoning methods. This could be at a more general level (e.g., prescribing principles about different treatment being necessary to correct for historical disparities) or more context-specific level (e.g., that only occupations with existing representation disparities for marginalized groups should have affirmative action). And finally, for correlation tasks which may naturally appear in real-world use (e.g., creative story-telling about characters), we follow the recommendation of prior works to design human-centered interventions Yee et al. (2021); Bennett et al. (2021). In other applications, e.g., when translating the gender-neutral phrase “o bir doktor” in Turkish into either “he is a doctor” or “she is a doctor” in English,666https://blog.google/products/translate/reducing-gender-bias-google-translate/ this has looked like providing multiple options to the user. This intervention pushes the user to make explicit decisions, rather than implicitly prioritizing certain choices over others and making potentially biased decisions.
@@
First, a key limitation of our benchmark suite is that, like most benchmark suites, it primarily measures upstream performance with uncertain predictability of downstream performance Wagstaff (2012). While our benchmarks are more downstream than correlation evaluations, they may still be distinct from a specific application, e.g., writing a recommendation letter Wan et al. (2023), autocompleting emails. Our intention is that performance on our benchmark suite is indicative of performance on other downstream applications. However for these reasons, our benchmark is intended to be understood on a relative rather than absolute scale. In App. E.5 we do an analysis of the within-benchmark correlation to try and better understand this, with the idea that if performance on our benchmarks have correlation with themselves, then performance on them is likely to correlate with other downstream applications which may warrant difference awareness. This is related to the problem that our benchmark is composed of multiple choice questions, which have been shown to not necessarily correlate with other kinds of uses Röttger et al. (2024a); Tam et al. (2024). However, there are benefits to multiple choice questions. Beyond being easier to analyze, there is a lower computational cost compared to open-ended responses. To ensure the difference unawareness we observe from our benchmark suite does not only manifest in the multiple choice setting, we sanity check open-ended versions of our questions on Google Gemini and Anthropic Claude chat interfaces. Both of the examples in our introduction are from this open-ended setting, and we include details in App. E.1.
@@
Second, our benchmark suite is not exhaustive in scope. Four of our eight benchmarks are explicitly grounded in the United States context, and while the other four may generalize to other contexts, are likely still based on Western norms and values Sambasivan et al. (2021). Examples of scenarios not included in the coverage of our benchmark suite include reclamations of slurs (i.e., members of certain identity groups using words that would otherwise be deemed inappropriate) Jeshion (2020), what composes a hate crime, additional diversity initiatives beyond affirmative action in the occupation setting, medical reasons to treat people from different demographic groups differently.
@@
Finally, there are a set of limitations resulting from our particular usage of demographic group. For one, we do not disaggregate. In other words, the scores combine outputs on questions asking about racial differences, gender differences, and more. While in many cases it can be important to disaggregate by demographic axis, our focus in this work is on demonstrating the erasure of difference awareness as an important concept for fairness. We encourage future work to explore whether and how difference awareness varies across demographic axes. Another key limitation based in demographic groups is the harm of group essentialization. In other words, that we may be reifying and legitimizing identities as rigid and innate. Furthermore, by treating identities as discrete categories, we alienate individuals who are outside of categories, for instance, non-binary and Multiracial. We hope this initial exploration into the relevance of group categories can lead to broader discussions about how the full spectrum of human experiences and identities can be recognized.
Overall, our benchmark suite offers the ability to measure two dimensions of fairness: difference awareness and contextual awareness. However, this does not capture aspects such as wrongful discrimination. Instead, we focused on complementing existing bias benchmark suites that already ensure groups are treated the same. Thus, results on our suite alone may not capture the bias of a model that discriminates unfairly between groups.
@@ Acknowledgments
SK acknowledges support by NSF 2046795 and 2205329, IES R305C240046, ARPA-H, the MacArthur Foundation, Schmidt Sciences, OpenAI, and Stanford HAI. We are appreciative of feedback from RegLab and STAIR Lab.
@@ Appendix B Literature Review
In Tbl. 4 we list the 37 existing language model fairness benchmark papers that we review, and how we categorized them for Tbl. 1 in the main text. Notably, this leaves out many canonical fairness measures that may be familiar to the reader, e.g., equal opportunity Hardt et al. (2016), because those are focused on the classification task rather than the generative one—even if evaluated on multiple choice responses. Certain benchmarks like HarmBench were left out because there was insufficient information from the paper to determine whether a difference unaware model would do well on them.
@@ Appendix E Other empirical results
The ten models we run our experiments on are Llama-3.1 8b and 70b Grattafiori et al. (2024), Mistral-0.3 7b Jiang et al. (2023), Mistral NeMo 12b Mistral AI (2024), Gemma-2 9b and 27b Team et al. (2024), GPT-4o regular and mini OpenAI et al. (2024), Claude-3.5 Sonnet and Haiku Anthropic (2024). For the Claude and GPT series we used the API, and for the other models we used Hugging Face. Here we describe the results of a series of additional experiments that we run.
@@ E.1 Introduction Examples
In Sec. 1 we include a few anecdotal examples of difference unawareness. These came from the chat interfaces of Claude and Gemini in January 2025. These were conducted in order to sanity check that we could naturally solicit examples of difference unawareness even without the constraints of forced multiple choice responses Röttger et al. (2024a); Tam et al. (2024). The full prompts and responses are included in Fig. 5.